I need help on question 5 and 6 and explaining it. Thank you
Annual Salary in 5 Years Quantiles -Summary Statistics Δ. Confidence Intervals H S . . 1000% maximum 150 Mean 74.816667 Parameter Estimate Lower CI Upper CI 1-Alpha 74.81667 70.0473 79.58604 0.950 34.25 Std Err Mean 2.3834993 Std Dev 18.46251 15.64945 22.51802 0.950 18.462506 Mean 99.5% 97.5% 900% 75.0% 500% 250% 100% 2.5% 0.5% 50 Std Dev 100 Upper 95% Mean 79.586038 725 N 57.1 70047296 60 83.75 Lower 95% Mean quartile median quartie 60 49.525 49 49 40 60 80 100 120 140 160 00% minimum
media%2F70f%2F70f0e897-4374-4cfc-9cf7-30
Answer #1

Q 5) 95% COnfidence Interval

\overline{x} \pm t\times s /\sqrt{n} \Rightarrow 74.82\pm 2*\frac{18.46}{\sqrt{60}}

= 74.82\pm 4.77 \Rightarrow (70.05, 79.59)

Interpreattion: With 95% Confidence ,we can say that the annual salary in 5 years is between 70.05$ and 79.59$